Show that the line y = x + 2 is a tangent to the parabola y = x[sup]2[/sup]
1
Expert's answer
2010-10-21T12:12:53-0400
Let's find the points of intersection of two graphs: x2 = x +2 x2 - x -2 = 0 D = 1 + 8 = 9; x1 = (1 + 3)/2 = 2; x2 = (1 - 3) / 2 = -1;The line intersects the parabola at two points (2,4) and (-1, 1) and cannot be tangent to it.